### v2

# Predictable Compression Failures: Order Sensitivity and Information Budgeting for Evidence-Grounded Binary Adjudication

Abstract
Transformers used for evidence-grounded question answering with binary adjudication (e.g., support/refute or yes/no) can be highly sensitive to the order in which exchangeable evidence is presented, producing dispersion across permutations and unreliable attempted answers (“hallucinations” under a Bernoulli predicate). We treat evidence order as a nuisance variable and show that next-token training minimizes expected conditional description length over orderings, which can be close to Bayes-optimal in expectation while deviating under any fixed ordering. We quantify this expectation–realization gap via a Quantified Martingale Violation (QMV) bound that predicts $O(\log n)$ growth in permutation dispersion under harmonic positional sensitivity. We then derive the Expectation-level Decompression Law (EDFL), relating expected information budget to achievable reliability for Bernoulli predicates, and use it to define Bits-to-Trust (B2T), Risk-of-Hallucination (RoH), and the Information Sufficiency Ratio (ISR) with a fixed ISR-gating rule for answer/abstain decisions under permutation mixtures. On 3,059 grounded items from a five-benchmark evidence-grounded QA suite (FEVER, HotpotQA, NQ-Open, PopQA, and Controls), we observe logarithmic dispersion and Jensen gains from uniform permutation mixtures. In a pre-specified held-out audit (528 items), an ISR=1=1 gate attains 0.0–0.7% hallucination with 20.6–27.9% abstention (95% CIs).
Evidence-grounded QA, order sensitivity, abstention
[…]
1 Introduction
Transformers with positional encodings can be strong in-context learners, yet their predictions can vary substantially under permutations of the same evidence (Falck et al., 2024; Liu et al., 2024). This is a practical problem in evidence-grounded QA: the evidence set is intended to be exchangeable (order should not matter), but the model’s probability of answering correctly (or refusing) can change when chunks are reordered. We study when such order sensitivity produces unreliable attempted answers (“hallucinations” under a binary adjudication predicate) and connect it to an analytically derived answer/abstain interface based on information budgets under permutation mixtures.
Our starting point is that next-token training minimizes cross-entropy on sequences, which corresponds to minimizing expected conditional description length over orderings. This reconciles two facts: models can be close to Bayes-optimal on average while deviating from permutation-invariant predictors for any fixed ordering. We formalize this expectation-versus-realization gap via a dispersion bound (QMV) and connect it to a deployment interface: when the evidence order induces high dispersion and information is insufficient, abstain rather than guess (Algorithm 1). The resulting ISR gate is fixed at 1 by theory and depends only on per-permutation Bernoulli probabilities (not post-hoc threshold tuning).
This paper is not a general hallucination benchmark: we focus on evidence-grounded binary adjudication (verifiable Bernoulli predicates and answer/abstain decisions) to isolate the roles of order sensitivity and information budgets.
Qwen2-7B Llama3.1 00.10.10.20.20.30.30.40.4Slope bbDispersion scaling
Qwen2-7B Llama3.1 $5\cdot 10^{-2}0$.10.1Jensen gapMixture gain (nats/token)
Figure 1: Order sensitivity and mixture gains (Experiment 1; N=3,059N=3{,}059, $m=16$ banded permutations). Left: estimated log-scaling slope $b$ in the empirical law $|R_{\pi}|\approx a+b\log n$ (95% CI as defined in Section 4.1; Table 2). Right: Jensen gap (nats/token) from uniform permutation mixtures, quantifying the cross-entropy improvement from averaging over evidence orderings.
1.1 Our Contributions
We make three contributions:
  • Quantified order sensitivity. We prove a Quantified Martingale Violation (QMV) bound controlling permutation-induced dispersion, with explicit constants and $O(\log n)$ scaling under harmonic positional sensitivity (Theorem 1 and Theorem 2).
  • EDFL as an actionable Bernoulli specialization. EDFL (Theorem 5) is a Bernoulli coarse-graining of a convexity + data-processing bound, yielding closed-form planners (B2T/RoH/ISR) and a fixed ISR=1{=}1 answer/abstain rule; we also derive a rare-event lower bound scaling like $\log(1/\bar{q})$.
  • A deployable answer/abstain rule. From EDFL we define Bits-to-Trust (B2T), Risk-of-Hallucination (RoH), and the Information Sufficiency Ratio (ISR), and give an ISR-gating procedure under permutation mixtures (Algorithm 1).
1.2 Empirical Validation
We validate the theory using observables not defined by the planners themselves (Appendix B). In particular:
  • Permutation mixtures improve log-loss and accuracy (Experiment 1), and dispersion follows $a+b\log n$ across depths (Figure 1).
  • Randomized dose-response identifies a causal effect of information on hallucination, holding prompt length fixed (Experiment 2; Figure 2).
  • A pre-specified held-out audit reports an out-of-sample operating point for ISR gating on the reliability–coverage plane without post-hoc threshold tuning (Experiment 3; Figure 3).
[…]
Figure 2: Causal dose-response of information (Experiment 2). Randomized study holding prompt length fixed at L=4L{=}4 chunks while varying support dose $d\in\{0,1,2,3\}$ (subset of Factuality Slice; Appendix H.4). We plot the net change from dose $0\to 3$ in hallucination rate against the induced decrease in the information-budget estimate ($\Delta\bar{\Delta}:=\bar{\Delta}_{d=0}-\bar{\Delta}_{d=3}\approx 3\times 0.375=1.125$ nats). The fitted slope (OLS) is 0.127 fewer hallucinations per additional nat (12.7-12.7 pp/nat); over the same shift, answer rate increases by 37.5pp and accuracy on attempts by 45.6pp.
7070727274747676787880808282848400.20.20.40.40.60.60.80.811Coverage (%)Hallucination (%)
Figure 3: Audit operating point (Experiment 3; $N=528$, $m=6$). Pre-specified held-out audit with ISR=1=1 and fixed seeds. Shaded region shows the 95% CI for coverage (1–abstention) and hallucination rate (Wilson intervals as in Section 4.1); midpoint shown for visualization.
2 Related Work
Order sensitivity and permutation invariance. Transformers can be highly sensitive to evidence order, including systematic “lost-in-the-middle” effects (Liu et al., 2024), and formal violations of permutation invariance under positional processing (Falck et al., 2024). We focus on quantifying the resulting dispersion with an explicit $O(\log n)$ bound (QMV) and connecting it to an operational deployment interface.
Bayesian/MDL views of in-context learning. Prior work connects transformers and in-context learning to Bayesian inference in various forms (Xie et al., 2022; Zhang et al., 2023) (see also (Bai et al., 2023; Reuter et al., 2024)). The compression–learning connection is classically captured by Minimum Description Length (MDL) (Grünwald, 2007), and recent work relates language modeling and compression explicitly (Delétang et al., 2023). Our contribution is to make the expectation vs. realization gap over evidence orderings explicit and measurable (QMV and Jensen gaps), rather than treating order-invariance as an implicit property of the predictor.
Hallucination control, calibration, and abstention. There is extensive work on detection and calibration signals, e.g. semantic entropy for detecting errors (Farquhar et al., 2024) and lower bounds on calibrated predictors tied to Good–Turing-like effects (Kalai and Vempala, 2024). Several recent papers frame LLM behavior through lossy-compression lenses (Su and others, 2024; Vasilatos and others, 2024). Relative to this literature, we focus on evidence-grounded Bernoulli adjudication and provide a deployment-oriented budget–reliability law (EDFL) that yields a non-tunable ISR=1=1 answer/abstain rule under permutation mixtures. Our answer/abstain interface is also related to selective classification (reject option) and risk–coverage tradeoffs (Chow, 1970; Geifman and El-Yaniv, 2017); our emphasis is on the Bernoulli coarse-graining plus permutation mixtures that yield an audited, fixed-threshold ISR boundary.
3 Theory: Order-Sensitive Information-Theoretic Guarantees
We formalize how positional processing creates systematic deviations from Bayesian behavior while maintaining average-case optimality.
3.1 Setup and Notation
Let $x$ denote an instance consisting of $n$ evidence chunks $X=(x_{1},\ldots,x_{n})$ and a target variable $Y$. Let $\pi$ be a permutation of chunk indices $\{1,\ldots,n\}$ and let $\Gamma_{\pi}$ reorder the chunks, $\Gamma_{\pi}(X)=(x_{\pi(1)},\ldots,x_{\pi(n)})$. Let $S_{\pi}(\cdot)$ denote the model’s predictive distribution over $Y$ given $\Gamma_{\pi}(X)$. We use $P$ to denote a target distribution over $Y$ (ground truth in supervised evaluation). In deployment, we instantiate $P$ with a reference distribution $P_{\mathrm{ref}}$ induced by the application’s specification for a predicate $g$ (gold labels in supervised evaluation; deterministic verification primitives such as database lookups, unit tests, or rules; human review; or a calibrated verifier). Our guarantees are stated relative to $P_{\mathrm{ref}}$.
Importantly, we do not need to construct a reference distribution over free-form outputs. For a binary adjudication predicate $g:\mathcal{Y}\to\{0,1\}$, define the induced (pushforward) distributions on $\{0,1\}$:
$P_{\mathrm{ref},g}(z)$ $=P_{\mathrm{ref}}(g(Y)=z),$
$S_{\pi,g}(z)$ $=S_{\pi}(g(Y)=z),\qquad z\in\{0,1\}.$
By data processing (coarse-graining cannot increase KL), $\operatorname{KL}(P_{\mathrm{ref}}\|S_{\pi})\geq\operatorname{KL}(P_{\mathrm{ref},g}\|S_{\pi,g})$, so the relevant reference object for answer/abstain decisions is the Bernoulli parameter $p_{\mathrm{ref}}:=P_{\mathrm{ref},g}(1)$. For a Bernoulli predicate $g:\mathcal{Y}\to\{0,1\}$ (e.g., correctness, answer-vs-refuse, constraint satisfaction), define $q_{\pi}(x):=S_{\pi}(g{=}1)$ and $\bar{q}(x):=\mathbb{E}_{\pi}[q_{\pi}(x)]$. We assume $P_{\mathrm{ref}}\ll S_{\pi}$ for all $\pi$ (absolute continuity), smoothing zero-probability tokens with $\varepsilon=10^{-9}$ in practice. We write $E_{\text{pair}}:=\mathbb{E}_{\pi,\pi^{\prime}}|q_{\pi}(x)-q_{\pi^{\prime}}(x)|$ for the expected pairwise absolute difference across independent permutations (with $\pi,\pi^{\prime}$ i.i.d. uniform over permutations). All expectations over $\pi$ are conditional on fixed instance $x$.
All information quantities (KL divergences, information budgets, Jensen gaps) are reported in nats unless otherwise noted.
3.2 Quantified Martingale Violations
Definition 1 (Permutation-induced residual).
For a Bernoulli predicate $g$ and position-aware predictions $q_{\pi}(x):=S_{\pi}(g{=}1)$, the permutation residual is $R_{\pi}(x):=q_{\pi}(x)-\bar{q}(x)$ where $\bar{q}(x):=\mathbb{E}_{\pi}[q_{\pi}(x)]$.
Assumption 1 (Local rank stability with bounded total variation).
For a fixed item $x$, define $f_{x}(\pi):=\operatorname{logit}(q_{\pi}(x))$ on $S_{n}$. For each chunk $i\in\{1,\dots,n\}$ and rank $t\in\{1,\dots,n{-}1\}$ define the coordinate-wise adjacent increment
$\Delta_{i,t}\;:=\;\sup_{\pi_{-i}}\Big|f_{x}\big(r_{i}{=}t{+}1,\pi_{-i}\big)-f_{x}\big(r_{i}{=}t,\pi_{-i}\big)\Big|,$
where $\pi_{-i}$ ranges over permutations of the remaining chunks and $r_{i}$ is the rank of chunk $i$. Let the coordinate total variation be $\mathrm{TV}_{i}(x):=\sum_{t=1}^{n-1}\Delta_{i,t}$ and suppose
$\sum_{i=1}^{n}\mathrm{TV}_{i}(x)\;\leq\;B(x)\;<\;\infty.$
Moreover, if there exist nonnegative coefficients $C_{i}$ and $\alpha>0$ such that $\Delta_{i,t}\leq C_{i}\,t^{-\alpha}$ and $\sum_{i}C_{i}=:C_{\mathrm{tot}}(x)<\infty$, we say the decay is $(\alpha,C_{\mathrm{tot}})$-regular.
Theorem 1 (Quantified Martingale Violation).
Under Assumption 1,
$\mathbb{E}_{\pi}\big|R_{\pi}(x)\big|\;\leq\;\mathbb{E}_{\pi,\pi^{\prime}}\big|q_{\pi}(x)-q_{\pi^{\prime}}(x)\big|\;\leq\;\frac{1}{4}\,\sum_{i=1}^{n}\mathrm{TV}_{i}(x).$
If, in addition, the decay is $(\alpha,C_{\mathrm{tot}})$-regular, then
$\mathbb{E}_{\pi}\big|R_{\pi}(x)\big|$ $\;\leq\;\frac{C_{\mathrm{tot}}(x)}{4}\times$
[…]
In words. QMV bounds permutation-induced dispersion by the cumulative sensitivity of the logit to adjacent rank changes. Under harmonic decay ($\alpha=1$), the bound grows like $\log n$, motivating the empirical $a+b\log n$ scaling test in Experiment 1.
Assumption 2 (First-order positional sensitivity).
There exist nonnegative content weights $w_{1},\ldots,w_{n}$ with $\sum_{i}w_{i}=1$ and a potential $\psi$ that is $(\alpha,C)$-regular (i.e., $|\psi(r+1)-\psi(r)|\leq Cr^{-\alpha}$) such that:
$\operatorname{logit}(q_{\pi}(x))=a(x)+\sum_{i=1}^{n}w_{i}\psi(\text{pos}_{\pi}(i))$
Proposition 1 (First-order model $\Rightarrow$ local stability).
Suppose Assumption 2 holds with nonnegative weights $w_{i}$ summing to $1$ and $\psi$ $(\alpha,C)$-regular. Then Assumption 1 holds with $\Delta_{i,t}\leq w_{i}\,C\,t^{-\alpha}$ and hence $C_{\mathrm{tot}}(x)=C$.
Theorem 2 (Explicit log-scaling under first-order positional sensitivity).
Under Assumption 2 with harmonic decay ($\alpha=1$), the permutation-induced dispersion admits an explicit $O(\log n)$ upper bound with constants:
$\mathbb{E}_{\pi}|R_{\pi}(x)|\leq\frac{C}{4}(\log n-\frac{3}{2}+o(1))$
More generally: $\alpha<1\Rightarrow O(n^{1-\alpha})$; $\alpha=1\Rightarrow O(\log n)$; $\alpha>1\Rightarrow O(1)$ due to harmonic versus $p$-series convergence.
Proposition 2 (Assumption-free JS certificate).
Let $\bar{S}=\mathbb{E}_{\pi}S_{\pi}$. For any Bernoulli predicate $g$ with $q_{\pi}=S_{\pi}(g{=}1)$ and $\bar{q}=\bar{S}(g{=}1)$,
$\mathbb{E}_{\pi}\big|q_{\pi}-\bar{q}\big|\;\leq\;\mathbb{E}_{\pi}\operatorname{TV}(S_{\pi},\bar{S})\;\leq\;\sqrt{\tfrac{1}{2}\,\mathbb{E}_{\pi}\operatorname{KL}(S_{\pi}\;\|\;\bar{S})}.$
Note $\mathbb{E}_{\pi}\operatorname{KL}(S_{\pi}\|\bar{S})$ is the (generalized) Jensen-Shannon divergence (JSD) with uniform weights, hence the bound controls dispersion by JSD via Pinsker.
3.3 MDL Optimality Through Architectural Closure (conceptual)
How to read this subsection. The results below are representational: they formalize that finite permutation mixtures are realizable if one allows tied-weight multi-branch evaluation with an averaging head (i.e., ensembling “inside” the hypothesis class). This is a conceptual lens on expectation-versus-realization gaps and does not claim that a standard single-branch deployed transformer literally implements the multi-branch architecture or that SGD recovers the mixture without additional structure. In practice, expectation over orderings comes either from training-time exposure to randomized orderings or from inference-time permutation mixtures.
Theorem 3 (Permutation-mixture realizability with averaging head).
Fix any base model $p_{\theta}(y|x)$ in the model family and any finite set of permutations $\Pi$. Consider an ensemble-within-the-network that:
  1. (i)
    Applies the same parameters $\theta$ to $x$ along each branch after permuting inputs by $\Gamma_{\pi}$ for every $\pi\in\Pi$
  2. (ii)
    Outputs per-branch distributions $p_{\theta}(y|\Gamma_{\pi}(x))$
  3. (iii)
    Averages the distributions in probability space with equal weights
Then the composite network implements:
$q_{\theta,\Pi}(y|x)=\frac{1}{|\Pi|}\sum_{\pi\in\Pi}p_{\theta}(y|\Gamma_{\pi}(x))$
If the model class includes such tied-weight multi-branch compositions with a linear averaging head, it is closed under finite permutation mixtures. As $|\Pi|\to\infty$ along i.i.d. draws from the uniform measure over permutations, $q_{\theta,\Pi}\Rightarrow\bar{p}_{\theta}(y|x)=\mathbb{E}_{\pi}[p_{\theta}(y|\Gamma_{\pi}(x))]$ almost surely.
Proof.
The only operation beyond the base model is a convex combination in probability space. Since each branch outputs a normalized distribution, the average is also a normalized distribution and is implementable by a fixed linear layer that sums corresponding probabilities across branches with weights $1/|\Pi|$. Tied weights ensure the branches are copies of $p_{\theta}$ evaluated at permuted inputs, so the composite realizes the desired mixture. By the strong law of large numbers, the empirical average over i.i.d. permutations converges pointwise to the permutation expectation. ∎
Theorem 4 (MDL Optimality in Expectation).
Under the architectural closure of Theorem 3, the permutation-mixture family can achieve information-theoretic optimality in expectation over orderings. Write $P_{T}:=P(\cdot|T(X))$ and $p_{\theta,\pi}:=p_{\theta}(\cdot|\Gamma_{\pi}(X))$. The risk decomposes as:
$\mathbb{E}_{X,Y,\pi}\big[-\log p_{\theta,\pi}(Y)\big]$ $=H(Y|T)$
$\quad+\mathbb{E}_{X,\pi}\operatorname{KL}\!\big(P_{T}\,\|\,p_{\theta,\pi}\big)$
The gap between expectation and realization is captured by the positional Jensen penalty:
$\mathfrak{J}_{\Gamma}(P,\theta):=\mathbb{E}_{\pi}[\operatorname{KL}(P\|p_{\theta,\pi})]-\operatorname{KL}(P\|\bar{p}_{\theta})\geq 0$
Over the convex hull of permutation mixtures realized by the architecture, the I-projection onto the exchangeable target $P(\cdot|T(X))$ yields:
$\inf_{\theta}\mathbb{E}_{(X,Y)}\mathbb{E}_{\pi}\big[-\log p_{\theta,\pi}(Y)\big]=H(Y|T)+o(1)$
Proof.
The risk decomposition follows from the chain rule for KL divergence. Since Theorem 3 guarantees the model family contains $\bar{p}_{\theta}$, the convex hull of permutation mixtures is realizable. The I-projection of $P(\cdot|T(X))$ onto this convex hull minimizes the KL divergence, and when the target is realizable (i.e., exchangeable and in the convex hull), the residual term vanishes, yielding MDL optimality in expectation. ∎
3.4 The Expectation-level Decompression Law
Binary adjudication is the natural unit of analysis as EDFL provides closed-form bounds for Bernoulli events and production systems make binary answer/abstain decisions (we evaluate via Bernoulli predicates $g$, details in Appendix B).
Theorem 5 (Expectation-level Decompression Law (EDFL)).
For any event $\mathcal{A}$ with prior mass $\bar{q}=\bar{S}(\mathcal{A})$ and posterior mass $p=P(\mathcal{A})$, the expected information budget satisfies:
$\bar{\Delta}:=\mathbb{E}_{\pi}[\operatorname{KL}(P\|S_{\pi})]\geq\operatorname{KL}(\operatorname{Ber}(p)\|\operatorname{Ber}(\bar{q}))$
with equality when $P$ is the $I$-projection of $\bar{S}$ onto $\{Q:Q(\mathcal{A})=p\}$. Equality holds for the exponentially tilted distribution $P^{\star}(y)\propto\bar{S}(y)e^{\lambda g(y)}$ with $\lambda$ chosen so $P^{\star}(\mathcal{A})=p$ (standard $I$-projection).
Remark 1 (Bernoulli instantiation).
Fix a predicate $g$ and consider the induced distributions $P_{\mathrm{ref},g}$ and $S_{\pi,g}$ on $\{0,1\}$. Applying EDFL on this coarse-grained space yields the information budget
$\bar{\Delta}_{g}$ $=\mathbb{E}_{\pi}[\operatorname{KL}(P_{\mathrm{ref},g}\|S_{\pi,g})]$
$=\mathbb{E}_{\pi}\big[\operatorname{KL}(\operatorname{Ber}(p_{\mathrm{ref}})\|\operatorname{Ber}(q_{\pi}))\big],$
which depends only on Bernoulli probabilities. By data processing, $\mathbb{E}_{\pi}[\operatorname{KL}(P_{\mathrm{ref}}\|S_{\pi})]\geq\bar{\Delta}_{g}$, so working in Bernoulli space is conservative when the downstream decision is binary.
Novelty note. EDFL itself is an immediate consequence of convexity of KL and data processing (Appendix A.4). We highlight it because the Bernoulli coarse-graining yields a closed-form, deployment-facing rule (B2T/RoH/ISR) with a fixed ISR=1=1 threshold under permutation mixtures.
In words. EDFL lower-bounds the expected information budget needed to shift an event’s probability from prior mass $\bar{q}$ to posterior mass $p$. For rare events, the required budget grows like $\log(1/\bar{q})$ (Corollary 1), which motivates B2T as a minimum budget for target reliability and the ISR gate as a deployment-time rule.
Corollary 1 (Compression Failure for Rare Events).
For fixed $p=1-\varepsilon$ with $\varepsilon\in(0,\tfrac{1}{2}]$, when $\bar{q}\ll 1$, achieving reliability $p=1-\varepsilon$ requires:
$\bar{\Delta}\geq(1-\varepsilon)\log\frac{1}{\bar{q}}+O(\bar{q})$
As a uniform lower bound for all $\varepsilon\in(0,\tfrac{1}{2}]$:
$\bar{\Delta}\geq\frac{1}{2}\log\frac{1}{\bar{q}}-\log 2+O(\bar{q})$
Insufficient information can lead to compression failure manifesting as hallucination.
3.5 Operational Planners
EDFL uses $p$ to denote the (unknown) achieved success probability of an event under the reference distribution. In contrast, planning specifies a target success probability $p^{\star}=1-h^{\star}$, and B2T/ISR are computed relative to this target.
Box 1: Hallucination Prevention Metrics
$\text{B2T}(x;p^{\star})=\operatorname{KL}(\operatorname{Ber}(p^{\star})\|\operatorname{Ber}(q_{\text{lo}}(x)))$ $\text{RoH}(x)=1-p_{\max}(\bar{\Delta}(x),\bar{q}(x))$ $\text{ISR}(x)=\bar{\Delta}(x)/\operatorname{B2T}(x;p^{\star})$ Target: set $p^{\star}=1-h^{\star}$. Decision rule: ISR < 1 → abstain; ISR $\geq 1$ → answer
From EDFL, the abstain/answer rule uses a fixed ISR threshold of 1.0 determined analytically. Evidence permutations, clipping (B=6B{=}6), and seeds were set before scoring, ensuring boundary alignment is a falsifiable out-of-sample check. B2T, RoH, and ISR are derived from EDFL as operational planners for deployment-time decisions.
A worked numerical example translating EDFL into B2T/RoH/ISR and an answer/abstain decision appears in Appendix B.3.
Algorithm 1 ISR gating with permutation mixture (Bernoulli instantiation)
0: Prompt $x$, target hallucination rate $h^{\star}$ (so $p^{\star}=1-h^{\star}$), permutations $m$, clip $B$, predicate $g$, reference $p_{\mathrm{ref}}:=P_{\mathrm{ref}}(g(Y){=}1)$
1: Sample permutations $\{\pi_{k}\}_{k=1}^{m}$ of evidence; form inputs $\Gamma_{\pi_{k}}(x)$
2:for $k=1$ to $m$ do
3:  Query model $\to$ $q_{k}:=S(g(Y){=}1\mid\Gamma_{\pi_{k}}(x))$
4:  Compute budget term $u_{k}:=\operatorname{KL}(\operatorname{Ber}(p_{\mathrm{ref}})\parallel\operatorname{Ber}(q_{k}))$
5:end for
[…]
10:if $\mathrm{ISR}\geq 1$ then
11:  return Answer (generate with guardrails)
12:else
13:  return Abstain (or acquire information and re-evaluate)
14:end if
In our supervised experiments, $P_{\mathrm{ref}}$ is induced by gold labels and $p_{\mathrm{ref}}$ is a point mass (with $\varepsilon$-smoothing), so $u_{k}$ reduces to the usual negative log-likelihood under each permutation. In deployment, $P_{\mathrm{ref}}$ should be read as a specification oracle / verification primitive for the predicate $g$.
3.5.1 Deployment instantiations (operationalizing $P_{\mathrm{ref}}$)
Algorithm 1 requires a Bernoulli reference $p_{\mathrm{ref}}$ for the predicate $g$. This is not available pre-answer in many open-world grounded QA settings. The intended deployments are those where the predicate is verifiable either (i) before generating long-form content (adjudication or tool-backed decisions), or (ii) after drafting but before release (post-hoc release filtering/retry). Three concrete instantiations are:
  1. (i)
    Binary adjudication with a deterministic oracle (pre-answer). Let the system’s output be a binary adjudication $Y\in\{0,1\}$ (e.g., approve/deny, support/refute) and let a rules engine or database lookup provide $y_{\mathrm{ref}}(x)\in\{0,1\}$ at decision time. Set $P_{\mathrm{ref}}=\delta_{y_{\mathrm{ref}}(x)}$, choose $g(y)=y$, so $p_{\mathrm{ref}}=y_{\mathrm{ref}}(x)$, and compute $q_{k}=S(g{=}1\mid\Gamma_{\pi_{k}}(x))=S(Y{=}1\mid\Gamma_{\pi_{k}}(x))$ from label-token probabilities (as in Experiment 1).
  2. (ii)
    Multiple-choice / discrete-tool answers (pre-answer). Let the answer be one of $K$ discrete options and let a tool/KB return the correct option $y_{\mathrm{ref}}(x)\in\{1,\dots,K\}$. Define $g(y)=\mathbb{1}_{\{y=y_{\mathrm{ref}}(x)\}}$, so $p_{\mathrm{ref}}=1$, and compute $q_{k}=S(g{=}1\mid\Gamma_{\pi_{k}}(x))=\sum_{y:\,g(y)=1}S(Y{=}y\mid\Gamma_{\pi_{k}}(x))$ via normalized option-token probabilities.
  3. (iii)
    Post-hoc verifiable generation (pre-release). Let $y$ be a drafted artifact (e.g., code, a structured query) and let a verifier $V(x,y)\in\{0,1\}$ implement the predicate (unit tests, constraint checkers). Apply Algorithm 1 as a release gate: first draft $y$, then set $p_{\mathrm{ref}}:=V(x,y)$ and compute $q_{k}$ by asking a predictor/verifier model for $\Pr[V(x,y)=1\mid\Gamma_{\pi_{k}}(x)]$ under evidence permutations. If the verifier returns a calibrated probability rather than a hard label, use that value as $p_{\mathrm{ref}}$; guarantees are then conditional on verifier calibration.
4 Experiments
Experimental setup. We evaluate on a five-benchmark evidence-grounded QA suite: four established real-world benchmarks (FEVER, HotpotQA, NQ-Open, PopQA) plus a Controls benchmark (insufficient-evidence and recency-trap items). We standardize these benchmarks into a common evidence-grounded format that we call Factuality Slice; the combined suite contains 3,059 items with controlled evidence chunks and hard negatives (Appendix H). Experiment 1 studies permutation dispersion and mixture gains under binary labels; Experiments 2–3 evaluate answer/abstain decisions.
Meta-benchmark assembly. We standardize each benchmark into the same evidence-grounded schema (question, evidence chunks with provenance, support spans, and hard negatives), with fixed chunking/capping and BM25-based negative mining. The suite is reproducible by construction: we pin SHA-256 hashes for all raw downloads and use fixed-seed train/val/test splits; Appendix H documents inclusion criteria, licensing, and build details.
Metrics. We report coverage (1–abstention), accuracy on attempts (task-correct among answered items), and hallucination rate (fraction of answered items with $g(Y)=0$) for the binary adjudication predicate $g$ (Appendix B); these need not coincide, so we report both. For the held-out audit we also report boundary alignment: the fraction of items for which the ISR decision agrees with adjudication (answer when g=1g{=}1, abstain when g=0g{=}0). Uncertainty, compute, and reproducibility conventions are summarized in Section 4.1; dataset build and licensing details are in Appendix H.
4.1 Experimental reporting and reproducibility
Uncertainty.
For proportions (coverage/abstention, hallucination rate, accuracy on attempts, boundary alignment) we report 95% Wilson score intervals over items. For the dispersion slopes $b$ in Table 2, we fit OLS to the 58 per-$n$ mean residuals and report the standard 95% confidence interval for the slope coefficient. Unless explicitly shown, other quantities are point estimates.
Compute/resources.
We report algorithmic compute via the total number of model forward passes (one forward pass per prompt/permutation to obtain next-token label probabilities). Evidence chunks are capped at 48 tokens, so the evidence portion ranges from $3\times 48$ to $60\times 48$ tokens (plus question/instruction text).
Table 1: Compute summary (algorithmic). Total forward passes are computed as $\text{items}\times m\times\text{models}$, where $m$ is the number of evaluated evidence permutations per item.
Experiment Models Items $\times\,m$ Total fwd passes
Exp. 1 (disp./mix.) 2: Qwen2-7B Llama-3.1-8B $3{,}059\times 16$ 97,88897{,}888
Exp. 3 (audit) Gemma-2-9B FT $528\times 6$ 3,1683{,}168
Reproduction path.
To preserve double-blind review, URLs for code and processed datasets are withheld in the main paper; we provide full methodological detail in the paper/appendix and will release the dataset builder and experiment scripts at camera-ready. Fine-tuning details for the Gemma-2-9B audit model (Experiment 3): LoRA with r=32r{=}32, $\alpha{=}64$, dropout 0.05 targeting attention and MLP projection modules; BF16 with FlashAttention-2; max sequence length 2048; gradient accumulation 64 with per-device batch size 1; 1 epoch; learning rate $1.5\times 10^{-5}$; weight decay 0.1; cosine schedule with 0.03 warmup ratio; seed 42.
4.2 Experiment 1: Dispersion scaling and mixture gains
Question.
How large is permutation-induced dispersion under exchangeable evidence, and do uniform permutation mixtures yield the expected Jensen gains (and near-optimal mixtures) in practice?
4.2.1 Dispersion scaling (QMV)
Design.
We conduct large-scale dispersion studies on 3,059 binary classification items from all splits spanning $n\in[3,60]$ chunks (58 distinct $n$ values). Each item contains 48-token-capped evidence chunks with binary gold labels. We test Qwen2-7B-Instruct and Llama-3.1-8B-Instruct, both with 4-bit NF4 quantization. For each item we draw $m=16$ unique banded permutations (6 bands, shuffle within), compute predictions $q_{\pi}(x)=P_{\pi}(\text{"1"})/(P_{\pi}(\text{"1"})+P_{\pi}(\text{"0"}))$ via renormalized label token probabilities, and form the uniform mixture $\bar{q}(x)=\frac{1}{m}\sum_{k=1}^{m}q_{\pi_{k}}(x)$.
Metric.
We measure mean absolute residual $|R_{\pi}|=\mathbb{E}_{\pi}|q_{\pi}(x)-\bar{q}(x)|$ and fit the empirical law $|R_{\pi}|\approx a+b\log n$ (Figure 1, Table 2).
Results.
Mean absolute residual $|R_{\pi}|$ follows $a+b\log n$ across both models, confirming Theorem 1. Qwen2-7B shows stronger positional sensitivity than Llama-3.1-8B, reflecting architectural differences. Mean absolute residuals remain approximately 69% of $E_{\text{pair}}$ at $n=60$ for both models.
Table 2: Summary of permutation dispersion and mixture effects (Experiment 1). N=3,059N=3{,}059 items; banded permutations with $m=16$ per item and $n\in[3,60]$ chunks (48-token cap). Bracketed intervals for $b$ are 95% CIs as defined in Section 4.1.
Metric Qwen2-7B Llama-3.1-8B
Dispersion Slope $b$ (vs. $\log n$) 0.377 (0.319, 0.435) 0.147 (0.109, 0.184)
$R^{2}$ 0.742 0.515
Jensen Gap (nats/token) 0.1041 0.00982
Mixture Optimality Gap $10^{-4}$ $\leq 5.3\times 10^{-5}$
4.2.2 Mixture gains and near-optimality
We fix a canonical continuation $Y$ once per item (chosen from the identity ordering by argmax over the two label probabilities), score that same $Y$ under $m=16$ banded permutations of the evidence ($n\in[3,60]$), and compare single-permutation vs. mixture behavior. For each item we compute the Jensen gap
$\underbrace{\mathbb{E}_{k}[-\log S_{k}(Y)]}_{\text{mean single-perm CE}}\;-\;\underbrace{-\log\mathbb{E}_{k}[S_{k}(Y)]}_{\text{uniform mixture CE}}\;\geq 0,$
where $S_{k}(Y)$ is the model’s probability of $Y$ under permutation $k$ (two-label normalization). We also learn a global convex mixture over permutations (shared weights per $n$) by exponentiated-gradient and report its cross-entropy.
  • Qwen-2-7B-Instruct (3,059 items; $m=16$; $n\in[3,60]$). Mean Jensen gap = 0.1041 nats/token (uniform mixture strictly improves over single permutations). The uniform mixture is essentially optimal: the mixture-optimality gap (uniform vs. globally optimized mixture) is $10^{-4}$ nats/token. The mean per-$n$ Jensen gaps are positive across the entire range $n=3\ldots 60$ (Appendix C).
  • Llama-3.1-8B-Instruct (same setup). Mean Jensen gap = 0.00982 nats/token; the mixture-optimality gap is $\leq 5.3\times 10^{-5}$ nats/token. Again, means are positive for every $n$.
Together these results confirm the Jensen inequality at the item level (positive gaps) and show that uniform permutation mixtures achieve virtually the same cross-entropy as the globally optimized mixture, i.e., near-MDL optimality within the permutation family. (All numbers are in nats/token; full per-$n$ tables and weight vectors appear in Appendix C.)
4.3 Experiment 2: Causal dose-response of information
Question.
Does increasing information content causally reduce hallucination when prompt length is held fixed?
Design.
To identify the effect of information content (not prompt length), we run randomized experiments holding the prompt length fixed at $L=4$ chunks while varying the dose $d\in\{0,1,2,3\}$ of support chunks versus non-support chunks.
Identification.
Dose randomization shifts the information budget $\bar{\Delta}$ while holding length constant. Appendix D reports robustness checks (including IV using $d$ as an instrument for $\bar{\Delta}$).
Results.
From dose 0 to 3, answer rate increases by 37.5 percentage points (pp), accuracy by 45.6pp, and hallucination decreases by 17.6pp. The information-budget estimate decreases by 0.375 nats per additional support chunk (Spearman $\rho=-0.80$, p<0.001), and the OLS slope is 0.127 fewer hallucinations per additional nat of decrease. A separate Llama-3.1-8B validation yields $\beta\approx 0.110$ hallucination reduction per nat. Figure 2 summarizes these dose-response relationships. We use symmetric clipping for stability; min-clipping provides a provable lower bound (Appendix A.5).
Interpretation.
Because dose shifts $\bar{\Delta}$ while holding length constant, the negative dose-response supports the EDFL mechanism: information sufficiency improves with supportive evidence and reduces hallucination risk under Bernoulli adjudication (Theorem 5).
4.4 Experiment 3: Pre-specified held-out audit
Question.
Does the fixed ISR=1=1 boundary generalize out-of-sample as a reliability–coverage operating point without threshold tuning?
Design.
We evaluate a Gemma-2-9B model fine-tuned on the four non-control benchmarks in Factuality Slice (FEVER, HotpotQA, NQ-Open, PopQA) on a 528-item held-out audit drawn from the full five-benchmark suite (including Controls). We pre-specify the permutation seeds $\{0,1,\ldots,5\}$ (so $m=6$), preserve role markers, and use symmetric clipping with $B=6$ nats. Boundary alignment is the fraction of items for which the ISR gate’s decision agrees with adjudication (answer when g=1g{=}1, abstain when g=0g{=}0). In supervised evaluation, $P_{\mathrm{ref}}$ is induced by the gold label, so the budget reduces to $\bar{\Delta}=\frac{1}{m}\sum_{k=1}^{m}\operatorname{clip}(-\log q_{k},B)$ where $q_{k}=S_{k}(g{=}1)$ is the model probability assigned to the correct outcome.
Results.
Boundary alignment is 96.2% [94.3, 97.5], hallucination rate is 0.0–0.7%, abstention is 24.1% [20.6, 27.9], and accuracy on attempts is 80.5% [76.8, 83.8]. The mean Jensen gap is $\hat{\mathfrak{J}}_{\Gamma}=0.82$ nats. Figure 3 reports the corresponding reliability–coverage operating point.
Interpretation and robustness.
[…]
5 Synthesis
The three experiments provide complementary validation, each mapping directly to our theoretical contributions. Experiment 1 validates Theorem 1 (QMV) through observed $O(\log n)$ scaling and Theorem 4 (MDL optimality) through strictly positive Jensen gaps. Experiment 2 validates Theorem 5 (EDFL) by showing causal dose-response of hallucination to information budgets. Experiment 3 demonstrates the operational planners derived from EDFL work in practice with an ISR=1=1 operating point separating safe answering from abstention.
6 Discussion
Implications.
In grounded QA, evidence order acts as a nuisance variable: different permutations induce measurable dispersion in predictions (Experiment 1; Figure 1). QMV explains why dispersion can grow with context depth under mild positional sensitivity (Theorem 1 and Theorem 2), while permutation mixtures recover log-loss improvements consistent with Jensen’s inequality. EDFL links reliability to information budgets for Bernoulli adjudication (Theorem 5); randomized dose-response supports that increasing information reduces hallucination (Experiment 2; Figure 2).
Operational takeaway.
ISR gating yields a concrete abstention interface on the reliability–coverage plane (Experiment 3; Figure 3): when ISR is low, abstain or acquire information; when ISR is high, attempt an answer.
6.1 Limitations
  • Binary adjudication focus. EDFL is sharpest for Bernoulli predicates; extensions to multi-class or structured generation require careful predicate design.
  • Reference for the predicate. In evaluation, $P_{\mathrm{ref}}$ comes from gold labels; in deployment, it comes from the system’s verification primitive or a calibrated verifier. Guarantees are relative to $P_{\mathrm{ref}}$ and must account for verifier error.
  • Compute overhead. Permutation mixtures require multiple model calls ($m$), trading latency for reliability.
  • Permutation scheme dependence. Banded vs uniform permutations (and chunking choices) can change dispersion constants; we partially probe sensitivity in Appendix E.
  • Scope of evidence grounding. Results are established on an evidence-grounded QA dataset; generalization to open-ended generation is not claimed.
6.2 Broader impact and ethics
Potential positive impacts. ISR-style gating can make abstention a first-class, auditable decision in evidence-grounded QA. Potential negative impacts. The same techniques increase compute (more model calls) and could be misused outside their assumptions; we emphasize scope limits (Section 6.1) and treat the verifier as part of the system that must be audited.
7 Conclusion
We connect order sensitivity in evidence-grounded QA to information budgets and a deployment-time abstention interface, yielding an ISR=1{=}1 gate with low hallucination.
Appendix A Appendix A: Proofs and Technical Details
A.1 A.1 Auxiliary Lemmas
Lemma 1 (Logistic Lipschitz).
Let $\sigma(t)=1/(1+e^{-t})$ and let $q=\sigma(u)$, $q^{\prime}=\sigma(v)$. Then
$|q-q^{\prime}|\leq\tfrac{1}{4}|u-v|.$
Proof.
By the mean value theorem, $|\sigma(u)-\sigma(v)|=|\sigma^{\prime}(\xi)||u-v|$ for some $\xi$ between $u$ and $v$. Since $\sigma^{\prime}(t)=\sigma(t)(1-\sigma(t))\leq 1/4$ for all $t$, the result follows. ∎
Lemma 2 (Regular potentials bound).
If $\psi$ is $(\alpha,C)$-regular, i.e., $|\psi(r{+}1)-\psi(r)|\leq Cr^{-\alpha}$ for $r\geq 1$, then for any integers $1\leq u,v\leq n$,
$|\psi(u)-\psi(v)|\leq\begin{cases}C\,H_{|u-v|},&\alpha=1,\\[2.0pt] \displaystyle\frac{C}{1-\alpha}\Big(|u-v|^{1-\alpha}-1\Big),&\alpha\in(0,1),\\[6.0pt] \displaystyle C\,\sum_{t=1}^{|u-v|}(t)^{-\alpha}\leq C\,\zeta(\alpha),&\alpha>1,\end{cases}$
where $H_{m}=\sum_{t=1}^{m}1/t$ and $\zeta$ is Riemann’s zeta function.
Lemma 3 (Harmonic-distance identity).
Let $U,V$ be i.i.d. uniform on $\{1,\dots,n\}$ and D=|UV|D=|U-V|. Then
$\mathbb{E}[H_{D}]\;=\;H_{n}-\frac{3}{2}+O\!\left(\frac{1}{n}\right).$
A.2 A.2 Proof of Theorem 2 (Explicit log-scaling under first-order positional sensitivity)
Proof of Theorem 2.
Write $q_{\pi}(x)=\sigma\!\big(a(x)+\sum_{i=1}^{n}w_{i}\psi(\mathrm{pos}_{\pi}(i))\big)$ under Assumption 2. For i.i.d. permutations $\pi,\pi^{\prime}$, by symmetrization,
$\mathbb{E}_{\pi}|R_{\pi}(x)|\;=\;\mathbb{E}_{\pi}\big|q_{\pi}(x)-\bar{q}(x)\big|\;\leq\;\mathbb{E}_{\pi,\pi^{\prime}}\big|q_{\pi}(x)-q_{\pi^{\prime}}(x)\big|.$
By Lemma 1,
$|q_{\pi}-q_{\pi^{\prime}}|\leq\tfrac{1}{4}\left|\sum_{i=1}^{n}w_{i}\big(\psi(\mathrm{pos}_{\pi}(i))-\psi(\mathrm{pos}_{\pi^{\prime}}(i))\big)\right|\leq\tfrac{1}{4}\sum_{i=1}^{n}w_{i}\big|\psi(\mathrm{pos}_{\pi}(i))-\psi(\mathrm{pos}_{\pi^{\prime}}(i))\big|.$
Taking expectation and using $\sum_{i}w_{i}=1$, exchangeability gives
$\mathbb{E}_{\pi,\pi^{\prime}}|q_{\pi}-q_{\pi^{\prime}}|\;\leq\;\tfrac{1}{4}\sum_{i=1}^{n}w_{i}\mathbb{E}\big|\psi(U)-\psi(V)\big|=\tfrac{1}{4}\,\mathbb{E}\big|\psi(U)-\psi(V)\big|,$
where $U,V$ are i.i.d. uniform on $\{1,\dots,n\}$. By Lemma 2, for $\alpha=1$, we have $\mathbb{E}|\psi(U)-\psi(V)|\leq C\cdot\mathbb{E}[H_{D}]$ where D=|UV|D=|U-V|. By Lemma 3, $\mathbb{E}[H_{D}]=H_{n}-\frac{3}{2}+o(1)=\log n-\frac{3}{2}+o(1)$. Thus:
$\mathbb{E}_{\pi}|R_{\pi}(x)|\;\leq\;\frac{C}{4}\,(\log n-\tfrac{3}{2}+o(1)).$
A.3 A.3 MDL Optimality and Positional Jensen Penalty
Proposition 3 (Risk decomposition and Jensen penalty).
For any $\theta$ and log-loss $\ell$:
$\mathbb{E}_{X,Y,\pi}[-\log p_{\theta}(Y|\Gamma_{\pi}(X))]=H(Y|T)+\mathbb{E}_{X,\pi}\operatorname{KL}\!\big(P(\cdot|T(X))\|p_{\theta}(\cdot|\Gamma_{\pi}(X))\big),$
where the Jensen penalty $\mathfrak{J}_{\Gamma}(X,\theta)=\mathbb{E}_{X,\pi}\operatorname{KL}(P\|p_{\theta}(\cdot|\Gamma_{\pi}))-\mathbb{E}_{X}\operatorname{KL}(P\|\bar{p}_{\theta})\geq 0$.
Theorem 6 (MDL Optimality via I-projection).
Under the architectural closure of Theorem 3, the convex hull of permutation mixtures contains all uniform mixtures $\bar{p}_{\theta}$. The I-projection onto the exchangeable target yields:
$\mathbb{E}[-\log q^{\dagger}(Y|X)]=H(Y|T)+\inf_{q\in\overline{\mathcal{Q}}}\mathbb{E}[\operatorname{KL}(P(\cdot|T)\|q)]$
The second term vanishes if the convex hull contains the exchangeable target.
A.4 A.4 EDFL and Corollary
Proof of Theorem 5.
Convexity of $Q\mapsto\operatorname{KL}(P\|Q)$ gives $\mathbb{E}_{\pi}\operatorname{KL}(P\|S_{\pi})\geq\operatorname{KL}(P\|\mathbb{E}_{\pi}S_{\pi})=\operatorname{KL}(P\|\bar{S})$. For the Bernoulli bound, apply data processing to $g:\mathcal{Y}\to\{0,1\}$, $g(y)=\mathbb{1}_{\{y\in\mathcal{A}\}}$, yielding $\operatorname{KL}(P\|\bar{S})\geq\operatorname{KL}(\operatorname{Ber}(p)\|\operatorname{Ber}(\bar{q}))$. Equality holds iff $g$ is sufficient for distinguishing $P$ from $\bar{S}$. ∎
Corollary 2 (Rare-event lower bound).
Fix $p=1-\varepsilon$ with $\varepsilon\in(0,\tfrac{1}{2}]$ and suppose $\bar{q}\in(0,1)$. As $\bar{q}\downarrow 0$:
$\operatorname{KL}(\operatorname{Ber}(p)\|\operatorname{Ber}(\bar{q}))\;\sim\;(1-\varepsilon)\log\frac{1}{\bar{q}}+O(\bar{q}).$
As a uniform lower bound for all $\varepsilon\in(0,\tfrac{1}{2}]$:
$\operatorname{KL}(\operatorname{Ber}(p)\|\operatorname{Ber}(\bar{q}))\;\geq\;\frac{1}{2}\log\frac{1}{\bar{q}}-\log 2+O(\bar{q}).$
A.5 A.5 Clipped Information-Budget Estimators
Let $U_{\pi}(y):=\log\frac{P(y)}{S_{\pi}(y)}$ and define:
$\widehat{\Delta}_{B}(y)=\frac{1}{m}\sum_{\pi=1}^{m}\mathrm{clip}(U_{\pi}(y),B)$
Proposition 4 (Lower-than-KL via min-clipping).
For any $B>0$ and any $P\ll S_{\pi}$:
$\mathbb{E}_{y\sim P}\big[\mathrm{clip}_{\min}(U_{\pi}(y),B)\big]\;\leq\;\operatorname{KL}(P\|S_{\pi})$
Consequently, $\mathbb{E}_{\pi}\mathbb{E}_{P}[\mathrm{clip}_{\min}(U_{\pi},B)]\leq\bar{\Delta}$.
A.6 A.6 Conditional-Complexity Interpretation
By the coding theorem, $K_{U}(y|\Gamma_{\pi}(x))=L_{\theta}(y|\Gamma_{\pi}(x))+O(1)$ in expectation. Therefore:
$\inf_{\theta}\mathbb{E}_{X,Y,\pi}[L_{\theta}(Y|\Gamma_{\pi}(X))]\;\approx\;\inf_{\theta}\mathbb{E}_{X,\pi}[K_{U}(Y|\Gamma_{\pi}(X))]$
By Theorem 4, this equals $H(Y|T)+o(1)$. Thus transformers minimize $\mathbb{E}_{\pi}[K_{U}(Y|\Gamma_{\pi}(X))]$, making them Bayesian in expectation over orderings, not in realization. All comparisons fix the universal machine; additive $O(1)$ constants cancel in differences and infima.
Appendix B Appendix B: Extended Methodology
B.1 B.1 Non-circular Validation Principle
We validate theory using observables not defined by our metrics: ground-truth likelihood/accuracy under permutation mixtures, the log-scaling of permutation-induced dispersion, and randomized dose-response of hallucination to information budgets. The planners we introduce (B2T, RoH, ISR) are decision rules, not validation targets; their threshold (ISR=1.0=1.0) is fixed ex ante by the derivation and is not tuned on evaluation data.
B.2 B.2 Binary Adjudication Rationale
We use Bernoulli outcomes because they are the natural unit for both theory and deployment. For Bernoulli events, EDFL reduces to a closed-form KL bound, yielding exact bits-to-trust thresholds; production systems decide to answer or abstain before emitting long-form content; any structured generation admits a verifiable predicate $g(y)\!\in\!\{0,1\}$ (factuality, constraint satisfaction, unit tests, rubric pass), so EDFL applies directly to $\Pr[g(Y)=1]$; and binary adjudication avoids grading/decoding confounds, isolating the information budget that drives hallucination risk.
B.3 B.3 Worked Example: From EDFL to a Decision
Setup. Target hallucination rate $h^{\star}=0.05$ (so $p^{\star}=0.95$). Compute B2T for conservative priors $q_{\text{lo}}$: $\mathrm{B2T}(p^{\star}=0.95;\,q_{\text{lo}})=\mathrm{KL}(\mathrm{Ber}(p^{\star})\,\|\,\mathrm{Ber}(q_{\text{lo}})).$ Numerics (nats): $\begin{array}[]{c|ccc}q_{\text{lo}}&0.02&0.10&0.30\\ \hline\cr\mathrm{B2T}&3.519&1.994&0.963\end{array}$ Budget $\bar{\Delta}$ to reliability. Given $\bar{q}=0.10$, the maximum achievable success at budget $\bar{\Delta}$ solves $\mathrm{KL}(\mathrm{Ber}(p)\,\|\,\mathrm{Ber}(0.10))\leq\bar{\Delta}$. Selected budgets $\to$ $p_{\max}$: $\begin{array}[]{c|cccc}\bar{\Delta}\ (\text{nats})&0.5&1.0&2.0&3.0\\ \hline\cr p_{\max}&0.495&0.689&0.951&\approx 1.000\end{array}$ ISR gate. If $\bar{q}=0.10$ and measured $\bar{\Delta}=2.0$, then $\mathrm{ISR}=\bar{\Delta}/\mathrm{B2T}=2.0/1.994\approx 1.00\Rightarrowanswer$. If instead $q_{\text{lo}}=0.02$, then $\mathrm{B2T}=3.519$ and the same budget gives $\mathrm{ISR}\approx 0.57\Rightarrowabstain$ or acquire info.
Appendix C Appendix C: Self-Consistency Detailed Results
C.1 Jensen gaps by depth ($n$)
Table 3: Self-Consistency (Jensen gaps). Positive (or zero) for all depths; zeros occur only at small $n\leq 6$ because banded permutations are degenerate there.
Qwen2-7B Llama-3.1-8B
Global mean Jensen gap (nats/token) 0.1041 0.00982
Per-$n$ range of mean gaps (nats/token) [0.0000, 0.2923] [0.0000, 0.03774]
% of $n$ with strictly >0>0 mean gap 54/58 54/58
Items (total) 3,059 3,059
C.2 Near-optimality of uniform mixtures
Table 4: Uniform permutation mixtures are essentially optimal. Cross-entropy (nats/token) for the self-consistency label $Y$, comparing (i) uniform mixture, (ii) globally optimized mixture (one convex weight vector shared per $n$), and (iii) oracle best single permutation (non-deployable lower bound).
Model Uniform CE Optimized CE Improvement Oracle single CE
(nats/token) (nats/token) (nats/token) (nats/token)
Qwen2-7B 0.2666 $\approx 0.2666$ $10^{-4}$ 0.3707
Llama-3.1-8B 0.28349 0.28344 $\leq 5.3\times 10^{-5}$ 0.29331
Appendix D Appendix D: Dose-Response Identification
Randomization and identification.
[…]
Appendix E Appendix E: Audit Parameter Sensitivity
Table 5: Parameter Sensitivity Analysis
Parameter Range Boundary Alignment Jensen Gap
Permutations (m) 3 94.7% 0.79
6 (default) 96.2% 0.82
12 97.1% 0.83
Clipping (B) 4 nats 95.1% 0.76
6 nats (default) 96.2% 0.82
8 nats 97.5% 0.84
Results remain robust across reasonable parameter ranges. The conditional independence audit shows MI($\mathcal{A}$, $\pi$) = 0.0032 nats (permutation test p=0.71).
Appendix F Appendix F: Rare Event Analysis Tables
Table 6: Information Sufficiency Determines Abstention
$\bar{q}$ $\bar{\Delta}$ (nats) $\operatorname{B2T}_{p^{\star}=0.95}$ (nats) ISR Decision
0.000 0.83 5.29 0.16 Refuse
0.042 1.91 3.78 0.51 Refuse
0.167 2.64 2.48 1.06 Answer
0.333 2.74 1.61 1.70 Answer
0.500 2.81 0.98 2.87 Answer
0.667 2.85 0.51 5.59 Answer
0.833 2.89 0.20 14.45 Answer
1.000 2.95 0.00 $\infty$ Answer
Appendix G Appendix G: Non-circularity and Threats to Validity
Our central empirical results (mixture gains on ground-truth labels, the $\log n$ dispersion law, and randomized dose-response) do not reference B2T/RoH/ISR. The ISR boundary report (96.2%) is a secondary boundary-alignment check with a pre-specified threshold (no tuning). Notably, 3.8% misalignment remains, which would be impossible under a tautological metric. Hence the main claims neither depend on, nor are validated by, the planners themselves.
Appendix H Appendix H: Factuality Slice Dataset Documentation
H.1 H.1 Meta-benchmark Overview
We evaluate on Factuality Slice, a meta-benchmark assembled from five evidence-grounded QA benchmarks: four established benchmarks (FEVER, HotpotQA, NQ-Open, PopQA) plus Controls. We standardize each benchmark into a common evidence-grounded format (question, evidence chunks, support spans, and hard negatives) to enable controlled permutation tests under a shared protocol. The resulting suite contains 3,059 binary adjudication items with controlled evidence presentation, enabling precise measurement of permutation sensitivity and information sufficiency.
Table 7: Factuality Slice: five-benchmark meta-benchmark. Caps are per-source sampling limits before filtering/standardization; the final suite used in this paper contains 3,059 items.
Benchmark Cap Task type Evidence source Why included
FEVER 2,000 Fact verification (T/F) Wikipedia evidence Clean binary adjudication with gold support/refute evidence
HotpotQA 2,000 Multi-hop QA Wikipedia passages Order sensitivity under multi-hop evidence chains
NQ-Open 1,000 Open-domain QA Retrieved Wikipedia Retrieval-style grounding with diverse entities
PopQA 500 Long-tail QA Retrieved Wikipedia Stress-test rare entities (popularity score <50<50)
Controls 300+ NEI + recency Retrieved / synthetic Calibration under insufficient or outdated evidence
H.2 H.2 Benchmarks and Composition
Factuality Slice integrates four established QA benchmarks plus a Controls benchmark:
FEVER (Fact Verification): Up to 2,000 claims converted to binary true/false questions with Wikipedia evidence from a June 2017 snapshot. Each claim paired with gold supporting/refuting evidence sentences and topically-related distractors via BM25 retrieval.
HotpotQA (Multi-hop Reasoning): Up to 2,000 questions requiring evidence synthesis across multiple Wikipedia articles. Preserves multi-hop structure with explicit support spans marking reasoning chains.
NQ-Open (Open-domain QA): Up to 1,000 questions from Natural Questions with evidence retrieved using a Wikipedia proxy corpus. Answer-bearing sentences marked as support with BM25-selected hard negatives.
PopQA (Long-tail Entities): Up to 500 questions about rare entities (popularity score <50<50) testing performance on uncommon knowledge. Evidence retrieved from a Wikipedia proxy with weak supervision.
Controls: 300+ FEVER “NOT ENOUGH INFO” claims plus synthetic recency traps with outdated evidence, testing model calibration on insufficient or misleading information.
Inclusion criteria (brief).
  • FEVER: Include SUPPORTS/REFUTES claims with at least one evidence sentence; map to binary True/False questions. FEVER NEI claims are reserved for Controls.
  • HotpotQA: Include questions with annotated supporting_facts and preserve multi-hop support spans.
  • NQ-Open / PopQA: Include items with non-empty question and answer; retrieve evidence from a Wikipedia proxy corpus and mark answer-bearing sentences as support (weak supervision).
  • PopQA long-tail filter: Restrict to popularity score <50<50.
  • Controls: Include insufficient-evidence items (FEVER NEI) and recency traps with outdated evidence.
H.3 H.3 Dataset Construction Pipeline
Evidence Chunking: All evidence sentences capped at 48 tokens to ensure consistent granularity. Each chunk tagged with source document, sentence ID, and retrieval score.
BM25 Retrieval System: Inverted-index BM25 (k1=1.2, b=0.75) built over 200,000 Wikipedia sentences. Used for finding topically-related distractors that don’t contain answer.
Support Span Annotation: Gold evidence sentences marked as support spans. For retrieved evidence, sentences containing answer string marked as supportive.
Hard Negative Mining: Retrieve up to 120 candidate distractors per sample via BM25, excluding gold evidence; the final context is capped (below).
Context Capping: Cap each sample to at most 60 evidence chunks, preserving all support spans and filling remaining slots with distractors. Enables testing at context lengths $n\in[3,60]$.
Data Integrity: SHA-256 hashes pinned for all source files with verification on each build. Train/val/test splits (80/10/10) with deduplication by (question, answer) pairs.
H.4 H.4 Usage in Experiments
Experiment 1 (Permutation Dispersion): Used all 3,059 items to measure permutation-induced dispersion across n $\in$ [3, 60]. Banded permutations (6 bands, shuffle within) applied to evidence chunks. Binary classification on sufficiency enables clean measurement of position sensitivity.
Experiment 2 (Dose-Response): Subset of items with exactly 4 chunks used. Support chunks (containing answer) vs non-support chunks randomized while holding total length constant. Enables causal identification of information content effect.
Experiment 3 (Frontier Audit): 528 held-out items from all sources used for boundary alignment testing. Pre-specified permutation seeds 0, 1, …, 5 applied before evaluation.
H.5 H.5 Key Dataset Properties
Information Gradation: Natural variation in evidence strength from strong (exact answer present) to weak (answer inferable) to insufficient (control examples).
Multi-hop Preservation: 42% of HotpotQA samples retain multi-hop structure with 2 or more supporting evidence pieces.
Answer Coverage: 76% of samples have answer string appearing in context (excluding controls), enabling verification of extraction vs hallucination.
Evidence Dating: All evidence tagged with snapshot dates (primarily 2017) to identify temporal misalignment.
Reproducibility: Dataset construction pins hashes and uses fixed random seeds. To preserve double-blind review, code and processed dataset URLs are withheld in the main paper and will be released at camera-ready.
H.6 H.6 Licensing and Ethics
All source datasets used under permissive licenses: FEVER (CC-BY-SA 4.0), HotpotQA (CC-BY-SA 4.0), NQ-Open (CC-BY 4.0), PopQA (MIT). No personally identifiable information included. Questions focus on factual knowledge rather than subjective opinions.